DISSOCIATION DYNAMICS OF THE IBr^{−}(CO_{2}) VAN DER WAALS CLUSTER: A DIRECT VIEW OF SOLVENT-DRIVEN NON-ADIABATIC TRANSITIONS

\underline{LEONID SHEPS}, ELISA M. MILLER, ROBERT PARSON, and W. CARL LINEBERGER^{a}, JILA, Department of Chemistry and Biochemistry, University of Colorado at Boulder, Boulder, CO 80309; MATTHEW A. THOMPSON, US Naval Research Lab, Washington, DC 20375; SAMANTHA HOR-VATH and ANNE B. McCOY, Department of Chemistry, The Ohio State University, Columbus, OH 43210.

We present the results of several time-resolved pump-probe experiments that explore the photodissociation of a simple dihalide anion, IBr^{−}, clustered with a single CO_{2} solvent molecule. A 100-fs pump laser pulse (λ=800 or 400 nm) initiates the dissociation of IBr^{−} on either the A’^{2}Π_{1/2} or the higher-lying B^{2}Σ_{1/2}^{+} state of the anion. A second, time-delayed, tunable 100-fs laser pulse probes the subsequent dynamics by photoelectron spectroscopy. In the bare anion the A’ and B states correlate exclusively to the I^{−} or the Br^{−} photoproducts, respectively. However, the addition of a single solvent molecule induces charge switching in a subset of the excited molecules midway through the dissociation, resulting in mixed photoproduct distributions (both I^{−} and Br^{−}) from both states. In the case of dissociation on the A’ state (which correlates asymptotically to I^{−}), the CO_{2} solvent molecule produces charge switching to Br^{−} in about 3% of the products, even though it cannot induce potential curve crossings. Electronic structure calculations and MD simulations point to a subtle mechanism for these non-adiabatic charge hops, involving the effects of the CO_{2} bending vibration to bridge the energetic gap of the charge transfer. Conversely, for dissociation on the higher-lying B state (which correlates asymptotically to Br^{−}), the photoproduct ratio is about 60% Br^{−}:40% I^{−}, indicating large amounts of charge switching. In this case it is energetically possible for the CO_{2} molecule to cause the B state to cross with the nearby a’^{2}Π_{1/2} state, which correlates with I^{−}. The charge switching mechanism, which is likely to be more complex than for the A’ state, will be discussed in detail.

aWe gratefully acknowledge the funding from NSF and AFOSR